Use either the slope-intercept form or the point-slope form to find an equation of each line. Write each result in slope-intercept form, if possible. Passes through (-3,6) and (-1, -4)
Answer by mananth(16946)

(y-y1)/(y1-y2) = (x-x1)/(x1-x2)
(x1,y1) (-3,6) and (x2,y2)(-1, -4)
(y-6)/(6-(-4))=(x-(-3))/(-3-(-1))
(y-6)/10 =(x+3)/-2
-2(y-6)=10(x+3)
-2y+6=10x+30
-2y=10x+24
/-2
y=-5x-12